t2=1/4 and x2=1/9

Answer by checkley71(8403)

T^2=1/4
T=SQRT1/4
T=1/2
PROOF
1/2^2=1/4
.5*.5=.25
.25=.25
------------------------------------------
X^2=1/9
X=SQRT1/9
X=1/3
PROOF
1/3^2=1/9
1/3*1/3=1/9